### v2

# Electronic Structure of Epitaxial Films of the Bilayer Strontium Ruthenate: Sr3Ru2O7

Abstract
We report the first combined study of the low-energy electronic band structure of epitaxial Sr3Ru2O7 films using angle-resolved photoemission spectroscopy and density functional theory. The complete Fermi-surface topography of the near-Fermi-level bands is determined from in-situ ARPES measurements. To investigate the effects of substrate-induced strain on the band structure, Sr3Ru2O7 thin films are epitaxially grown on SrTiO3 and (LaAlO3)0.3(Sr2TaAlO6)0.7 substrates using molecular beam epitaxy. The combination of the measured Fermi-surfaces along with the theoretical interpretation, clearly show dramatic changes in the Fermi surface topologies that result from the underlying strain states of the films on the two substrates. We find that the Sr3Ru2O7 films prepared on SrTiO3 are tensile strained with tetragonal symmetry, whereas those grown on (LaAlO3)0.3(Sr2TaAlO6)0.7 are compressively strained with orthorhombic symmetry. Within below the Fermi level, we observe two flat bands along - in the orthorhombic phase and around in the tetragonal phase. These features could be favorable for van Hove singularities near the Fermi level, and highlight the emergence of magnetic instabilities in epitaxial Sr3Ru2O7 films.
Ruthenium oxides of the Ruddlesden-Popper (R-P) phases, Srn+1RunO3n+1 (), are a class of strongly correlated materials that have garnered significant attention because they feature several comparable interactions that compete to engender a variety of novel electronic and magnetic phenomena [33, 51, 46, 5, 4]. The vast collection of ground-state properties in this series includes unconventional superconductivity in Sr2RuO4 () [33, 51], the quantum critical metamagnetism and nematicity in Sr3Ru2O7 () [46, 48], anisotropic ferromagnetism and proposed orbital-dependent metamagnetism in Sr4Ru3O10 () [5, 40] and the spontaneous itinerant ferromagnetism in SrRuO3 () [4]. The richness of different phenomena in Sr-based layered ruthenates comes from a competition between local and itinerant electronic behaviors. This involves a strong interplay between charge, spin, orbital, and lattice degrees of freedom, which underlie the complex ground states and responses of these materials to external perturbations [61, 44]. All the above ground-state properties demonstrate that the Srn+1RunO3n+1 materials provide a fascinating platform for exploring diverse novel quantum phenomena in correlated ruthenate systems.
The focus of this paper is on the bilayer strontium ruthenate system Sr3Ru2O7. It is a quasi-two-dimensional material due to the strong interactions of the Ru-O layers within the bilayer in the plane, and the weaker interactions between adjacent bilayers along the c-axis [20]. The crystal symmetry of Sr3Ru2O7 is commonly described as tetragonal, with space group I4/mmm [see, Fig. 1(a)]. However, structural refinements based on neutron and x-ray diffraction have revealed an in-plane rotation of the RuO6 octahedra by approximately 7, which lowers the symmetry [19, 53]. As a result, some studies report an orthorhombic structure, assigning space groups such as Bbcb or Pban, depending upon the specific modeling and refinement conditions [19, 54, 53, 18, 23].
Sr3Ru2O7 is a strongly enhanced Pauli paramagnet in its ground state, but undergoes a field-induced metamagnetic transition at temperatures as low as 5 K, characterized by a rapid and superlinear increase in magnetization. This metamagnetic behavior is accompanied by pronounced anomalies in thermodynamic and transport properties, including divergences in the electrical resistivity and the electronic specific heat near critical magnetic field values [27, 16, 45, 50, 37]. These strong fluctuations have been widely discussed within phenomenological frameworks invoking a putative metamagnetic quantum critical end point [16, 36, 14, 45].
There have also been studies suggesting that the microscopic origin of the metamagnetism in Sr3Ru2O7 lies in the electronic band structure. For instance, theoretical works and band structure calculations proposed that the phase diagram of Sr3Ru2O7 could be understood considering a Stoner-like model, where the resulting Fermi surface (FS) pocket derived from the Ru orbitals exhibits a nearby van Hove singularity in the density of states (DOS) [28, 2, 52, 54]. The predicted FS pockets of Sr3Ru2O7 have been experimentally observed in angle resolved photoemission spectroscopy (ARPES) measurements performed on single crystals [55]. Upon careful characterization of the band dispersions near the Fermi level (EF), it was found that the band presents saddle points and sharp peaks in the DOS, situations that are favorable for the occurrence of a van Hove singularity [55]. More importantly, the fact that these observations were made within an energy range consistent with the theoretical prediction of the Zeeman splitting for the metamagnetic transition [2] strengthens the idea that this emergent phenomenon in Sr3Ru2O7 can be directly traced to the detailed structure of its FS.
Consequently, the electronic band structure and the FS topology of Sr3Ru2O7 have been significantly investigated using both DFT calculations and ARPES [54, 26, 1]. So far, these studies have focused on Sr3Ru2O7 in its single crystal form. To gain more insight into the microscopic origin of the magnetic instability, it could be advantageous to explore the band structure and FS of thin film samples. Unlike single-crystals, thin films offer the opportunity to explore interfacial effects, and tailor the properties of the material based on strains induced by the substrates on which they are grown [34].
Any modification of the FS induced by strain, whether subtle or pronounced, is expected to influence the dynamics of magnetic-field-controlled instabilities and may give rise to new phases [27, 1, 3, 38]. Changes in the FS topology arise from alterations of the underlying crystal symmetry, allowing strain engineering to provide precise control over the electronic and magnetic interactions in bilayered Sr3Ru2O7 [62, 55, 11]. This tunability offers a powerful platform for exploring electronic and magnetic ordered phases, quantum criticality, and their relationship to the underlying band structure [58, 11, 38].
In particular, the strong sensitivity of Sr3Ru2O7 to small structural, electronic, or magnetic perturbations makes it an ideal system for examining how subtle modifications of lattice symmetry or orbital hybridization reshape the FS and stabilize or suppress competing ground states [62, 45]. Such control not only advances our understanding of correlated 4d electron systems but also opens pathways for engineering novel quantum phases through targeted modifications of the crystal environment. More broadly, strain engineering has proven effective for enhancing electronic and magnetic properties across materials, enabling improved photocatalytic and optoelectronic performance as well as controlled magnetic anisotropy in thin films [30, 56, 10]. However, electronic band-structure studies of epitaxial Sr3Ru2O7 remain limited by the challenges of controlling layered-ruthenate epitaxy and achieving phase-pure thin films [34, 8, 7].
Here, we report the first combined experimental and theoretical investigation of the electronic band structure and FS topology of Sr3Ru2O7 thin films. Using molecular beam epitaxy (MBE), we deposited Sr3Ru2O7 films on (LaAlO3)0.3(Sr2TaAlO6)0.7 (LSAT) and SrTiO3 (STO) single-crystal substrates. The reflection high-energy electron diffraction (RHEED) data acquired during growth indicate high-quality epitaxial Sr3Ru2O7 films; while the in-plane lattice parameters extracted from reciprocal space maps (RSMs) confirm coherent growth accompanied by compressive and tensile strains on LSAT and STO, respectively. To examine how these strain states modify the electronic structure, we performed ARPES measurements on the Sr3Ru2O7 films. We observe substantial strain-dependent transformations in the band dispersions. The measured FSs show that compressively strained films adopt an orthorhombic structure, whereas tensile-strained films exhibit a tetragonal symmetry. These observations indicate that the RuO6 octahedral rotation responsible for Brillouin zone reconstruction and band backfolding is suppressed in samples grown on STO. The DFT-calculated bands, obtained by incorporating the RSM-derived lattice parameters, are in general agreement with the ARPES experiments and reproduce the main features of the measured FSs. Within a few meV below EF, we observe two flat bands along - in the orthorhombic phase and around in the tetragonal phase. These bands suggest possible van Hove singularities near EF, a condition that favors magnetic instabilities in epitaxial Sr3Ru2O7 films.
To grow the Sr3Ru2O7 thin films, the LSAT and STO substrates were kept at the optimal temperature of 625 C, and the oxygen partial pressure in the growth chamber was maintained at 1.42 10-6 Torr. The procedure for the synthesis of the Sr3Ru2O7 films with high phase purity is described elsewhere [57]. For this study, LSAT with lattice parameter a = b = 3.868 Å [23, 17] and STO with lattice parameter a = b = 3.905 Å [23, 24] were chosen for controlled strain engineering, since their relatively small lattice mismatch with Sr3Ru2O7 (a = b = 3.890 Å) [19] yields respectively compressive and tensile strains in the film.
In-situ ARPES experiments were conducted at the temperature of 7 K, using a state-of-the art high-resolution Scienta DA-30L electron analyzer with an energy resolution of 5 meV. Figure 1(b) schematically illustrates the angle-resolved geometry of the ARPES experiment. For these measurements, 6 nm thick Sr3Ru2O7 samples were used. This was the optimal thickness ensuring high surface quality for the acquisition of high quality ARPES data. The theoretical investigations were carried out using pseudopotential, plane-wave implementation of DFT as described in Ref. [40, 13].
The growth of the Sr3Ru2O7 thin films was monitored in real time using RHEED. The diffraction patterns of the films grown on STO and LSAT are shown in Fig. 1(c) and 1(d), respectively. These patterns indicate good film quality and the epitaxial growth. In particular, the presence of the very sharp diffraction spots attests to the two dimensional growth mode of the Sr3Ru2O7 films on both substrates, as well as to the flatness, the smoothness, and the perfect single crystalline structure of their surfaces [22, 43, 7].
The high crystalline quality of the Sr3Ru2O7 films are further demonstrated in the structural data of representative films, presented in Fig. 1(e) - (g). The enlargement of -2 scans around the (0010) film peaks and the (002) substrate planes underlines the presence of Kiessig fringes [Fig. 1(e)]. These interference fringes highlight the high structural quality of the surfaces and the well defined interfaces [35]. The c-axis lattice parameter values extracted from these scans are 20.792 Å and 20.512 Å, for the samples grown on LSAT and STO, respectively. Compared with the reported bulk value of 20.719 Å [19], these values are consistent with both the vertical elongation and compression of the RuO6 octahedra associated with the in-plane compressive and tensile strains induced in the films by the LSAT and STO substrates, respectively [60].
Figures 1(f) and 1(g) show the reciprocal space maps around the (1015) Bragg reflections of the Sr3Ru2O7 films prepared on STO and LSAT, respectively. These maps indicate coherent epitaxial growth and fully strained films on both substrates, consistent with previous reports [34, 41, 8]. For scans around the (0115) as well as the (015) reflections, slightly different in-plane lattice parameters were obtained, especially for the films grown on LSAT. This would suggest an in-plane orthorhombicity for these films. However, given the limited resolution of the lab-based XRD system used for this study, these differences may fall within experimental errors, and could more conclusively be confirmed with high-resolution synchrotron measurements [59, 31]. Nevertheless, as will be discussed below, the orthorhombicity was incorporated into the FS calculations of the Sr3Ru2O7 films grown on LSAT, yielding good agreement with the measured data.
Strain, interfacial coupling and film thickness have been shown to induce rotations of the RuO6 octahedra in Sr-based ruthenate films, leading to modifications in both the underlying crystal symmetry and functional properties of the films [15, 21, 59, 31, 32]. In the following paragraphs, we examine how compressive and tensile strains alter the electronic band structures of Sr3Ru2O7 films and analyze the extent to which these changes correlate with strain-driven structural phase transitions associated with octahedral rotations.
Figure 2 presents the experimental and calculated FSs of Sr3Ru2O7 thin films prepared on both substrates. For the 6 nm thick film grown on LSAT [Fig. 2(a)], reasonable agreement is found with those of the orthorhombic Sr3Ru2O7 single crystal [55], although not all the predicted surface pockets are resolved in our experimental data. The square and cross shaped hole-like FS pockets (, ) originating from the and orbitals are clearly identifiable. The electron pockets and at the M point, as well as the innermost sheet at the point, are only weakly resolved in the ARPES data. A broad, high-intensity feature spanning the region near the X point is also observed, which may correspond to the small FS pocket labeled in orthorhombic Sr3Ru2O7 [55, 47]. The limited visibility of these small pockets is expected in epitaxial thin films, where reduced out-of-plane coherence, substrate-induced disorder, and strain-driven octahedral distortions broaden the spectral features, particularly for low–spectral-weight or symmetry-sensitive bands [29, 49].
The calculated FS [Fig. 2(b)] reproduces the main features observed in the measurements, including Brillouin-zone folding due to the doubled periodicity, the structural distortion inferred from the RSM analysis (orthorhombicity), and the effects of spin–orbit coupling. The latter likely explains the emergence of the pocket, whose orbital character is primarily , which is typically unoccupied in the single-layer ruthenate, but exhibits substantial hybridization with the states.
In the case of the 6 nm thick Sr3Ru2O7 film deposited on STO [Fig. 2(c)], the measured FSs closely resemble those of Sr3Ru2O7 in the ideal tetragonal structure [54]. This corresponds to the situation whereby, the FS analogous to that of the monolayer Sr2RuO4, reconstructs due to the interaction of the RuO2 sheets within the bilayer and their subsequent splitting and hybridization. The FS consists of two -centered electron-like pockets [54, 47] and a broad feature at . Although the orbital character of the latter is not yet firmly established, the former are believed to arise from an even combination of and orbitals [54].
The DFT results [Fig. 2(d)] show striking agreement with the ARPES measurements. The calculations assume that the in-plane Ru–O geometry is identical in the top and bottom layers, indicating no relative in-plane rotation of the RuO6 octahedra. This agreement suggests that, in the epitaxial bilayer Sr3Ru2O7 films, the electronic structure is predominantly determined by the Ru–O bonding geometry within each layer, while interlayer coupling and the thin-film constraints imposed by the substrate preserve the in-plane octahedral symmetry [12]. The absence of relative octahedral rotation or tilt implies minimal structural distortion, allowing the DFT calculations to accurately reproduce the measured FS features. Importantly, this highlights a thin-film effect: under epitaxial constraints, the octahedral network remains in a high-symmetry configuration, suppressing distortions or rotation patterns that could otherwise break symmetry, an effect not necessarily present in bulk crystals of similar composition.
The ARPES-measured FSs reveal substrate-dependent modifications that reflect the underlying strain states of the films. The tensile strain imposed by the STO substrate largely preserves tetragonal symmetry, whereas the compressive strain from the LSAT substrate reduces the symmetry and induces an orthorhombic phase. These structural differences, including the resulting fourfold rotation symmetries around , are well captured by the DFT calculations. Notably, while orthorhombic systems do not strictly possess fourfold symmetry, our observations remain consistent with previous ARPES and DFT studies on Sr3Ru2O7 single crystals [55, 47], indicating that the bilayer and substrate constraints dominate the observed electronic structure. Another important consideration is the low temperature at which the ARPES measurements are performed, as both LSAT and STO substrates are known to undergo symmetry changes that could, in principle, influence the film structure in addition to strain effects [6, 25]. However, the presence of orthorhombic and tetragonal symmetries at room temperature can be directly inferred from RSM measurements, reducing the likelihood that the observed structural changes arise from temperature-driven substrate transitions or interfacial coupling effects. To further corroborate the strain-driven structural distortions inferred from ARPES and DFT, the octahedral rotation patterns can be identified through half-integer reflections measured using synchrotron-based XRD, which are particularly sensitive to strain states and interfacial coupling [15, 21, 59, 31, 32]. Such measurements, however, are beyond the scope of the present study.
It is worth highlighting an interesting contrast between our measured and calculated FSs. This is the absence of the predicted innermost pocket in the orthorhombic phase, and the presence of an unexpected flat band feature at in the tetragonal phase. This is clearly laid out in Fig. 3, which shows the measured and calculated energy band dispersions along high symmetry directions. In the orthorhombic phase [Fig. 3(a) to 3(c)], the momentum distribution curves (MDCs) extracted at EF show no sign of dispersing feature around , while the band structure calculation shows two distinct centered bands crossing EF. These are bilayer splittings of the pocket [54]. They are unresolved in ARPES most likely because of their intrinsically low spectral weight, which in thin films can be further suppressed by disorder, reduced coherence, and matrix-element effect [55]. In the tetragonal phase on the other hand [Fig. 3(d) to 3(f)], the MDCs show a flat band near EF centered at , but no band dispersion crossing EF is seen in the calculation at . Instead, four nearly degenerate bands are observed well below EF. These are bonding–anti bonding splittings of the and states associated with the interbilayer interaction.
Note that the flat spectral weight character of the feature around 0.6 – 0.9 Å-1 along X in the orthorhombic phase [Fig. 3(c)] and that of the band at in the tetragonal phase [Fig. 3(e)] is consistent with the low-energy flat band reported in Ref. [42]. This is also confirmed through visualization of their edges [see supplemental information]. Owing to the proximity of these bands to EF (15 meV below EF), it could happen that nearby van Hove singularities occur and subsequently drive a metamagnetic instability in the Sr3Ru2O7 films when tuned by an applied magnetic field. This hypothesis, on the other hand, would require further validation through transport and specific-heat measurements, as both probes can reveal signatures of enhanced scattering and increased features in the DOS in the vicinity of the metamagnetic transition [45, 50, 2, 63].
Even with an energy resolution of 5 meV, resolving the finest band features in Sr3Ru2O7 thin films remains experimentally demanding. As discussed above, this behavior is consistent with modest quasiparticle linewidth broadening that can arise from surface roughness and other near-surface disorder inherent to complex oxide thin films [9]. While these effects currently limit a direct experimental identification of all FS pockets observed in bulk Sr3Ru2O7 single crystals [55], the overall agreement between ARPES and DFT demonstrates that the essential low-energy electronic structure is well captured in the thin-film geometry. Looking forward, high-resolution synchrotron-based ARPES measurements, combined with controlled light polarization (linear horizontal and linear vertical) [39], are expected to further enhance orbital selectivity and sensitivity to narrow features near EF, providing a promising pathway toward resolving finer details of the electronic structure in Sr3Ru2O7 thin films.
In conclusion, we have performed the first in-situ ARPES investigation of Sr3Ru2O7 thin films. For controlled strain engineering, we have prepared the Sr3Ru2O7 films on LSAT and STO to insure respectively compressive and tensile strains in the films. Upon ARPES measurements of the FSs, we found that the compressive strain promotes RuO6 octahedral rotations and leads to an orthorhombic symmetry, whereas the tensile strain retains a more symmetric, tetragonal-like structure. These observations are supported by the DFT calculations which implement the intrinsic structural characteristics of the grown samples. The ARPES measured band dispersions present narrow flat band features about 15 meV below EF. These bands could potentially be favorable for van Hove singularities and subsequently drive magnetic instabilities in the epitaxial Sr3Ru2O7 films. Our results underscore the strong sensitivity of the electronic structure of Sr3Ru2O7 to epitaxial strain and establish a foundation for future efforts to tune correlated phases in this bilayer ruthenate.
I Acknowledgements
P. N., S. S., and A. N. T. acknowledge funding from the College of Science and the Department of Physics & Astronomy at Louisiana State University. This work made use of the synthesis facility of the Platform for the Accelerated Realization, Analysis, and Discovery of Interface Materials (PARADIM), which is supported by the National Science Foundation under Cooperative Agreement No. DMR-2039380.
II Data Availability
The data that support the findings are available from the authors upon reasonable request.